THE MAGELLANIC CLOUDS NEWSLETTER

An electronic publication dedicated to the Magellanic Clouds, and astrophysical phenomena therein

No. 118 — 1 August 2012

http://www.astro.keele.ac.uk/MCnews

Editor: Jacco van Loon

Editorial

Dear Colleagues, It is my pleasure to present you the 118^{th} issue of the Magellanic Clouds Newsletter. Among the posted items are possible answers to questions about the origin of the sub-mm excess emission from the SMC, and about the origin of field O stars, as well as various other interesting new results. Why not consider attending the 30 Doradus workshop in Baltimore in September? The next issue is planned to be distributed on the 1^{st} of October 2012.

Editorially Yours, Jacco van Loon

Refereed Journal Papers

Stellar evolution of massive stars with a radiative α–ω dynamo

Adrian T. Potter^{1}, Shashikumar M. Chitre^{2} and Christopher A. Tout^{1}

1Institute of Astronomy, University of Cambridge, UK 2Centre for Excellence in Basic Sciences, University of Mumbai, India

Models of rotationally-driven dynamos in stellar radiative zones have suggested that magnetohydrodynamic transport of angular momentum and chemical composition can dominate over the otherwise purely hydrodynamic processes. A proper consideration of the interaction between rotation and magnetic fields is therefore essential. Previous studies have focused on a magnetic model where the magnetic field strength is derived as a function of the stellar structure and angular momentum distribution. We have adapted our one-dimensional stellar rotation code, rose, to model the poloidal and toroidal magnetic field strengths with a pair of time-dependent advection–diffusion equations coupled to the equations for the evolution of the angular momentum distribution and stellar structure. This produces a much more complete, though still reasonably simple, model for the magnetic field evolution. Our model reproduces well observed surface nitrogen enrichment of massive stars in the Large Magellanic Cloud. In particular it reproduces a population of slowly-rotating nitrogen-enriched stars that cannot be explained by rotational mixing alone alongside the traditional rotationlly-enriched stars. The model further predicts a strong mass-dependency for the dynamo-driven field. Above a threshold mass, the strength of the magnetic dynamo decreases abruptly and so we predict that more massive stars are much less likely to support a dynamo-driven field than less massive stars.

Accepted for publication in MNRAS Available from arXiv:1205.6477

The sub-mm and mm excess of the SMC: Magnetic dipole emission from magnetic nanoparticles?

B.T. Draine^{1} and Brandon Hensley^{1}

1Princeton University, USA

The Small Magellanic Cloud (SMC) has surprisingly strong sub-mm- and mm-wavelength emission that is inconsistent with standard dust models, including those with emission from spinning dust. Here we show that the emission from the SMC may be understood if the interstellar dust mixture includes magnetic nanoparticles, emitting magnetic dipole radiation resulting from thermal fluctuations in the magnetization. The magnetic grains can be metallic iron, magnetite (Fe_{3}O_{4}), or maghemite (γ-Fe_{2}O_{3}). The required mass of iron is consistent with elemental abundance constraints. The magnetic dipole emission is predicted to be polarized orthogonally to the normal electric dipole radiation if the nanoparticles are inclusions in larger grains. We speculate that other low-metallicity galaxies may also have a large fraction of the interstellar Fe in magnetic materials.

Submitted to ApJ Available from arXiv:1205.6810

Non-thermal excitation and ionization in supernovae

Chengdong Li^{1}, D. John Hillier^{1} and Luc Dessart^{2}

1Department of Physics and Astronomy & Pittsburgh Particle physics, Astrophysics, and Cosmology Center (PITT PACC), University of Pittsburgh, Pittsburgh, PA 15260, USA 2Laboratoire d’Astrophysique de Marseille, Université Aix-Marseille & CNRS, UMR 7326, 38 rue Frédéric Joliot–Curie, 13388 Marseille, France

We incorporate non-thermal excitation and ionization processes arising from non-thermal electrons that result from γ-ray energy deposition, into our radiative transfer code cmfgen. The non-thermal electron distribution is obtained by solving the Spencer–Fano equation using the procedure of Kozma & Fransson (1992). We applied the non-thermal calculations to the blue supergiant explosion model whose early evolution was studied in Dessart & Hillier (2010). Nonthermal processes generally increase excitation and ionization and decrease the temperature of the ejecta. We confirm that non-thermal processes are crucial for modeling the nebular spectra. Both optical H i and He i lines are significantly strengthened. While optical He i lines are not easily discerned in observational spectra due to severe blending with other lines, He i 2.058 μm provides an excellent opportunity to infer the influence of non-thermal processes. We also discuss the processes controlling the formation of the He i lines during the nebular epoch. Most lines of other species are only slightly affected. We also show that the inclusion of Fe i has substantial line-blanketing effects on the optical spectra. Our model spectra and synthetic light curves are compared to the observations of SN 1987A. The spectral evolution shows broad agreement with the observations, especially Hα. The uncertainties of the non-thermal solver are studied, and are expected to be small. With this new addition of non-thermal effects in cmfgen, we now treat all known important processes controlling the radiative transfer of a supernova ejecta, whatever the type and the epoch.

Accepted for publication in Monthly Notices of the Royal Astronomical Society Available from arXiv:1206.0215

Field O stars: formed in situ or as runaways?

V.V. Gvaramadze^{1}, C. Weidner^{2,3}, P. Kroupa^{4} and J. Pflamm-Altenburg^{4}

1Sternberg Astronomical Institute, Lomonosov Moscow State University, Russia 2Instituto de Astrofísica de Canarias, Spain 3Scottish Universities Physics Alliance (SUPA), School of Physics and Astronomy, University of St. Andrews, Scotland 4Argelander-Institut für Astronomie, Universita”at Bonn, Germany

A significant fraction of massive stars in the Milky Way and other galaxies are located far from star clusters and star-forming regions. It is known that some of these stars are runaways, i.e. possess high space velocities (determined through the proper motion and/or radial velocity measurements), and therefore most likely were formed in embedded clusters and then ejected into the field because of dynamical few-body interactions or binary-supernova explosions. However, there exists a group of field O stars whose runaway status is difficult to prove via direct proper motion measurements (e.g., in the Magellanic Clouds) or whose (measured) low space velocities and/or young ages appear to be incompatible with their large separation from known star clusters. The existence of this group led some authors to believe that field O stars can form in situ. Since the question of whether or not O stars can form in isolation is of crucial importance for star formation theory, it is important to thoroughly test candidates of such stars in order to improve theory. In this paper, we examine the runaway status of the best candidates for isolated formation of massive stars in the Milky Way and the Magellanic Clouds by searching for bow shocks around them, by using the new reduction of the Hipparcos data, and by searching for stellar systems from which they could originate within their lifetimes. We show that most of the known O stars thought to have formed in isolation are instead very likely runaways. We show also that the field must contain a population of O stars whose low space velocities and/or young ages are in apparent contradiction with the large separation of these stars from their parent clusters and/or the ages of these clusters. These stars (the descendants of runaway massive binaries) cannot be traced back to their parent clusters and therefore can be mistakenly considered as having formed in situ. We argue also that some field O stars could be detected in optical wavelengths only because they are runaways, while their cousins residing in the deeply embedded parent clusters might still remain totally obscured. The main conclusion of our study is that there is no significant evidence whatsoever in support of the in situ proposal on the origin of massive stars.

Accepted for publication in MNRAS Available from arXiv:1206.1596

Multi-frequency observations of SNR J0453−6829 in the LMC; A composite supernova remnant with a pulsar wind nebula

F. Haberl^{1}, M.D. Filipović^{2}, L.M. Bozzetto^{2}, E.J. Crawford^{2}, S.D. Points^{3}, W. Pietsch^{1}, A.Y. De Horta^{2}, N. Tothill^{2}, J.L. Payne^{2} and M. Sasaki^{4}

1MPE Garching, Germany 2University of Western Sydney, Australia 3CTIO, La Serena, Chile 4IAAT Tübingen, Germany

The Large Magellanic Cloud (LMC) is rich in supernova remnants (SNRs) which can be investigated in detail with radio, optical and X-ray observations. SNR J0453−6829 is an X-ray and radio-bright remnant in the LMC, within which previous studies revealed the presence of a pulsar wind nebula (PWN), making it one of the most interesting SNRs in the Local Group of galaxies. We study the emission of SNR J0453−6829 to improve our understanding of its morphology, spectrum, and thus the emission mechanisms in the shell and the PWN of the remnant. We obtained new radio data with the Australia Telescope Compact Array and analysed archival XMM–Newton observations of SNR J0453−6829. We studied the morphology of SNR J0453−6829 from radio, optical and X-ray images and investigated the energy spectra in the different parts of the remnant. Our radio results confirm that this LMC SNR hosts a typical PWN. The prominent central core of the PWN exhibits a radio spectral index α_{Core} = −0.04 ± 0.04, while in the rest of the SNR shell the spectral slope is somewhat steeper with α_{Shell} = −0.43 ± 0.01. We detect regions with a mean polarisation of P ∼ (12 ± 4)% at 6 cm and (9 ± 2)% at 3 cm. The full remnant is of roughly circular shape with dimensions of (31 ± 1) pc ×(29 ± 1) pc. The spectral analysis of the XMM–Newton EPIC and RGS spectra allowed us to derive physical parameters for the SNR. Somewhat depending on the spectral model, we obtain for the remnant a shock temperature of around 0.2 keV and estimate the dynamical age to 12000–15000 years. Using a Sedov model we further derive an electron density in the X-ray emitting material of 1.56 cm^{−3}, typical for LMC remnants, a large swept-up mass of 830 M_{⊙}, and an explosion energy of 7.6 × 10^{50} erg. These parameters indicate a well evolved SNR with an X-ray spectrum dominated by emission from the swept-up material.

Accepted for publication in Astronomy & Astrophysics Available from arXiv:1206.5679

Chemical abundances of metal-poor RR Lyrae stars in the Magellanic Clouds

Raoul Haschke^{1}, Eva K. Grebel^{1}, Anna Frebel^{2.3}, Sonia Duffau^{1}, Camilla Hansen^{4} and Andreas Koch^{4}

1Astronomisches Rechen-Institut, Zentrum für Astronomie der Universität Heidelberg, Mönchhofstraße 12–14, 69120 Heidelberg, Germany 2Harvard-Smithsonian Center for Astrophysics, 60 Garden St., MS-20, Cambridge, MA 02138, USA 3Massachusetts Institute of Technology, Kavli Institute for Astrophysics and Space Research, 77 Massachusetts Avenue, Cambridge, MA 02139, USA 4Landessternwarte, Zentrum für Astronomie der Universität Heidelberg, Königstuhl 12, 69117 Heidelberg, Germany

We present for the first time a detailed spectroscopic study of chemical element abundances of metal-poor RR Lyrae stars in the Large and Small Magellanic Cloud (LMC and SMC). Using the MagE échelle spectrograph at the 6.5m Magellan telescopes, we obtain medium resolution (R ∼ 2000–6000) spectra of six RR Lyrae stars in the LMC and three RR Lyrae stars in the SMC. These stars were chosen because their previously determined photometric metallicities were among the lowest metallicities found for stars belonging to the old populations in the Magellanic Clouds. We find the spectroscopic metallicities of these stars to be as low as [Fe/H]_{spec} = −2.7 dex, the lowest metallicity yet measured for any star in the Magellanic Clouds. We confirm that for metal-poor stars, the photometric metallicities from the Fourier decomposition of the lightcurves are systematically too high compared to their spectroscopic counterparts. However, for even more metal-poor stars below [Fe/H]_{phot} < −2.8 dex this trend is reversed and the spectroscopic metallicities are systematically higher than the photometric estimates. We are able to determine abundance ratios for ten chemical elements (Fe, Na, Mg, Al, Ca, Sc, Ti, Cr, Sr and Ba), which extend the abundance measurements of chemical elements for RR Lyrae stars in the Clouds beyond [Fe/H] for the first time. For the overall [α/Fe] ratio, we obtain an overabundance of 0.36 dex, which is in very good agreement with results from metal-poor stars in the Milky Way halo as well as from the metal-poor tail in dwarf spheroidal galaxies. Comparing the abundances with those of the stars in the Milky Way halo we find that the abundance ratios of stars of both populations are consistent with another. Therefore we conclude that from a chemical point of view early contributions from Magellanic-type galaxies to the formation of the Galactic halo as claimed in cosmological models are plausible.

Accepted for publication in Astronomical Journal Available from arXiv:1206.4999

Nitrogen chronology of massive main sequence stars

K. Köhler^{1}, M. Borzyszkowski^{1}, I. Brott^{2}, N. Langer^{1} and A. de Koter^{3,4}

1Argelander-Institut für Astronomie der Universität Bonn, Germany 2University of Vienna, Department of Astronomy, Austria 3Astronomical Institute, Utrecht University, The Netherlands 4Astronomical Institute Anton Pannekoek, University of Amsterdam, The Netherlands

Rotational mixing in massive main sequence stars is predicted to monotonically increase their surface nitrogen abundance with time. We use this effect to design a method for constraining the age and the inclination angle of massive main sequence stars, given their observed luminosity, effective temperature, projected rotational velocity and surface nitrogen abundance. This method relies on stellar evolution models for different metallicities, masses and rotation rates. We use the population synthesis code starmaker to show the range of applicability of our method. We apply this method to 79 early B-type main sequence stars near the LMC clusters NGC 2004 and N 11 and the SMC clusters NGC 330 and NGC 346. From all stars within the sample, 17 were found to be suitable for an age analysis. For ten of them, which are rapidly rotating stars without a strong nitrogen enhancement, it has been previously concluded that they did not evolve as rotationally mixed single stars. This is confirmed by our analysis, which flags the age of these objects as highly discrepant with their isochrone ages. For the other seven stars, their nitrogen and isochrone ages are found to agree within error bars, what validates our method. Constraints on the inclination angle have been derived for the other 62 stars,with the implication that the nitrogen abundances of the SMC stars, for which mostly only upper limits are known, fall on average significantly below those limits. Nitrogen chronology is found to be a new useful tool for testing stellar evolution and to constrain fundamental properties of massive main sequence stars. A web version of this tool is provided.

Accepted for publication in A&A Available from arXiv:1206.4146

Similarities in populations of star clusters

S. Michael Fall^{1} and Rupali Chandar^{2}

1Space Telescope Science Insitute, USA 2University of Toledo, USA

We compare the observed mass functions and age distributions of star clusters in six well-studied galaxies: the Milky Way, Magellanic Clouds, M 83, M 51, and Antennae. In combination, these distributions span wide ranges of mass and age: 10^{2}< ∼M/M_{⊙}< ∼10^{6} and 10^{6}< ∼τ/yr< ∼10^{9}. We confirm that the distributions are well represented by power laws: dN/dM ∝ M^{β} with β ≈ −1.9 and dN/dτ ∝ τ^{γ} with γ ≈ −0.8. The mass and age distributions are approximately independent of each other, ruling out simple models of mass-dependent disruption. As expected, there are minor differences among the exponents at a level close to the true uncertainties, ϵ_{β} ∼ ϵ_{γ} ∼ 0.1–0.2. However, the overwhelming impression is the similarity of the mass functions and age distributions of clusters in these different galaxies, including giant and dwarf, quiescent and interacting galaxies. This is an important empirical result, justifying terms such as ”universal” or ”quasi-universal.” We provide a partial theoretical explanation for these observations in terms of physical processes operating during the formation and disruption of the clusters, including star formation and feedback, subsequent stellar mass loss, and tidal interactions with passing molecular clouds. A full explanation will require additional information about the molecular clumps and star clusters in galaxies beyond the Milky Way.

Published in Astrophysical Journal Available from arXiv:1206.4237

Two compact H ii regions at the remote outskirts of the Magellanic Clouds

R. Selier^{1} and M. Heydari-Malayeri^{1}

1LERMA, Observatoire de Paris, France

The H ii regions LMC N 191 and SMC N 77 are among the outermost massive star-forming regions in the Magellanic Clouds. So far, few works have dealt with these objects despite their interesting characteristics. We aim at studying various physical properties of these objects regarding their morphology (in the optical and Spitzer IRAC wavelengths), ionized gas emission, nebular chemical abundances, exciting sources, stellar content, age, presence or absence of young stellar objects, etc. This study is based mainly on optical ESO NTT observations, both imaging and spectroscopy, coupled with other archive data, notably Spitzer images (IRAC 3.6, 4.5, 5.8, and 8.0 μm) and 2MASS observations. We show the presence of two compact H ii regions, a low-excitation blob (LEB) named LMC N 191A and a high-excitation blob (HEB) named SMC N 77A, and study their properties and those of their exciting massive stars as far as spectral type and mass are concerned. We also analyze the environmental stellar populations and determine their evolutionary stages. Based on Spitzer IRAC data, we characterize the YSO candidates detected in the direction of these regions. Massive star formation is going on in these young regions with protostars of mass about 10 and 20 M_{⊙} in the process of formation.

Accepted for publication in A&A Available from arXiv:1206.5939

GASKAP – The Galactic ASKAP Survey

John M. Dickey^{1}, Naomi McClure-Griffiths^{2}, Steven J. Gibson^{3}, José F. Gómez^{4}, Hiroshi Imai^{5}, Paul Jones^{6}, Snežana Stanimirović^{7}, Jacco Th. van Loon^{8}, Andrew Walsh^{9} and 49 additional authors

1University of Tasmania, Australia 2CSIRO, Australia 3Western Kentucky University, USA 4Instituto de Astrofísica de Andalucía, Spain 5Kagoshima University, Japan 6University of New South Wales, Australia 7University of Wisconsin, USA 8Keele University, UK 9James Cook University, Australia

A survey of the Milky Way disk and the Magellanic System at the wavelengths of the 21-cm atomic hydrogen (H i) line and three 18-cm lines of the OH molecule will be carried out with the Australian Square Kilometre Array Pathfinder telescope. The survey will study the distribution of H i emission and absorption with unprecedented angular and velocity resolution, as well as molecular line thermal emission, absorption, and maser lines. The area to be covered includes the Galactic plane (|b| < 10^{∘}) at all declinations south of δ = +40^{∘}, spanning longitudes 167^{∘} through 360^{∘} to 79^{∘} at b = 0^{∘}, plus the entire area of the Magellanic Stream and Clouds, a total of 13,020 square degrees. The brightness temperature sensitivity will be very good, typically σ_{T} ≃ 1 K at resolution 30′′ and 1 km s^{−1}. The survey has a wide spectrum of scientific goals, from studies of galaxy evolution to star formation, with particular contributions to understanding stellar wind kinematics, the thermal phases of the interstellar medium, the interaction between gas in the disk and halo, and the dynamical and thermal states of gas at various positions along the Magellanic Stream.

Accepted for publication in PASP Available from arXiv:1207.0891

Washington photometry of candidate star clusters in the Small Magellanic Cloud

Andrés E. Piatti^{1} and Eduardo Bica^{2}

1Instituto de Astronomía y Física del Espacio (IAFE), CC67, Suc 28, 1428, CABA, Argentina 2Universidade Federal do Rio Grande do Sul, Depto. de Astronomía CP 15051, Porto Alegre, 91500-970, Brazil

We present for the first time Washington CT1 photometry for 11 unstudied or poorly studied candidate star clusters. The selected objects are of small angular size, contain a handful of stars, and are projected towards the innermost regions of the Small Magellanic Cloud (SMC). The respective Colour–Magnitude Diagrams (CMDs) were cleaned from the unavoidable star field contamination by taking advantage of a procedure which makes use of variable size CMD cells. This method has shown to be able to eliminate stochastic effects in the cluster CMDs caused by the presence of isolated bright stars, as well as, to make a finer cleaning in the most populous CMD regions. Our results suggest that nearly 1/3 of the studied candidate star clusters would appear to be genuine physical systems. In this sense, the ages previously derived for some of them mostly reflect those of the composite stellar populations of the SMC field. Finally, we used the spatial distribution in the SMC of possible non-clusters to statistically decontaminate that of the SMC cluster system. We found that there is no clear difference between both expected and observed cluster spatial distributions, although it would become more important at a 2-σ level between a ∼ 0.^{∘}3 and 1.^{∘}2 (the semi-major axis of a ellipse parallel to the SMC bar and with b/a = 1/2), if the asterisms were increased up to 20%.

Accepted for publication in MNRAS Available from arXiv:1207.4769

The VMC Survey – VI. Quasars behind the Magellanic system

M.-R.L. Cioni^{1.2}, D. Kamath^{3}, S. Rubele^{4}, J.Th. van Loon^{5}, P.R. Wood^{3}, J.P. Emerson^{6}, B.K. Gibson^{7}, M.A.T. Groenewegen^{8}, V.D. Ivanov^{9}, B. Miszalski^{10,11} and V. Ripepi^{12}

1University Observatory Munich, Scheinerstraße 1, 81679 München, Germany 2University of Hertfordshire, Physics Astronomy and Mathematics, Hatfield AL10 9AB, United Kingdom 3Research School of Astronomy and Astrophysics, Mount Stromlo Observatory, Cotter Road, Weston Creek, ACT 2611, Australia 4INAF, Osservatorio Astronomico di Padova, Vicolo dell’Osservatorio 5, 35122 Padova, Italy 5Keele University, Lennard-Jones Laboratories, ST5 5BG, United Kingdom 6School of Physics and Astronomy, Queen Mary University of London, Mile End Road, London E1 4NS, United Kingdom 7Jeremiah Horrocks Institute, University of Central Lancshire, Preston PR1 2HE, United Kingdom 8Royal Observatory of Belgium, Ringlaan 3, 1180 Ukkel, Belgium 9European Southern Observatory, Av. Alonso de Córdoba 3107, Casilla 19, Santiago, Chile 10South African Astronomical Observatory, P.O. Box 9, Observatory, 7935, South Africa 11Southern African Large Telescope Foundation, P.O. Box 9, Observatory, 7935, South Africa 12INAF, Osservatorio Astornomico di Capodimonte, via Moiariello 16, 80131 Napoli, Italy

The present number and spatial distribution of confirmed quasi-stellar objects (QSOs) behind the Magellanic system is limited. This undermines their use as astrometric reference objects for detailed studies of proper motion, aimed at measuring rotation and stream motion in the galaxies, and to study the interstellar medium along the line of sight. We search for criteria to identify candidate QSOs using new near-infrared observations from the VISTA survey of the Magellanic Clouds system (VMC). The VMC survey provides photometry in the YJK_{s} wave bands and 12 epoch observations in the K_{s} wave band with unprecedented sensitivity and spatial resolution. The (Y–J) versus (J–K_{s}) diagram has been used to distinguish QSO candidates from foreground Milky Way stars and stars of the Magellanic Clouds. Then, the slope of variation in the K_{s} band has been used to identify a sample of high confidence candidates. These criteria were developed based on the properties of 117 known QSOs presently observed by the VMC survey where the K_{s} epochs span a time range up to 600 days. YJK_{s} magnitudes and K_{s} band light-curves of known QSOs behind the Magellanic system from present VMC data are presented. About 75% of known QSOs show a slope of variation in the K_{s} band > 10^{−4} mag day^{−1} and the shape of the K_{s} light-curve is in general irregular and without any clear periodicity. A method to identify QSOs based solely on the VMC data is proposed using YJK_{s} colours and Ks variability. QSO candidates were selected among VMC sources with photometric errors < 0.1 mag and quality flags = 0 in each wave band, Y < 19.32 mag, J < 19.09 mag, K_{s} < 18.04 mag and slope of variation in the K_{s} band > 10^{−4} mag day^{−1}. The number of QSO candidates found in tiles including the South Ecliptic Pole and the 30 Doradus regions is 22 and 26, respectively, with negligible contamination by young stellar objects, planetary nebulae, stars and normal galaxies. The high confidence in the nature of the selected objects is supported by recent studies of possible contaminants, but remains to be confirmed by spectroscopic follow up. By extrapolating the number of QSO candidates to the entire VMC survey area we expect to find about 1500 QSOs behind the LMC, 600 behind the SMC, 300 behind the Bridge and 50 behind the Stream areas. Without magnitude restrictions and allowing for de-blended objects, the VMC survey can find candidates > 1 mag fainter than those presented in this work. The VMC survey has the potential to reveal the remaining 85% of the QSOs behind the Magellanic system. Further, the K_{s} band light-curves can help support investigations of the mechanism responsible for the variations. The confirmed QSOs found from these candidates will provide a good astrometric reference grid for proper motion studies of the Magellanic system.

Submitted to Astronomy and Astrophysics Available from arXiv:1207.2492

On the excitation and formation of circumstellar fullerenes

J. Bernard-Salas^{1}, J. Cami^{2,3}, E. Peeters^{2,3}, A.P. Jones^{1}, E.R. Micelotta^{2} and M.A.T. Groenewegen^{4}

1Institut d’Astrophysique Spatiale, Université Paris-Sud 11, 91405 Orsay, France 2Department of Physics and Astronomy, The University of Western Ontario, London, ON N6A 3K7, Canada 3SETI Institute, 189 Bernardo Avenue, Suite 100, Mountain View, CA 94043, USA 4Royal Observatory of Belgium, Ringlaan 3, 1180 Brussels, Belgium

We compare and analyze the Spitzer mid-infrared spectrum of three fullerene-rich planetary nebulae in the Milky Way and the Magellanic Clouds; Tc 1, SMP SMC 16, and SMP LMC 56. The three planetary nebulae share many spectroscopic similarities. The strongest circumstellar emission bands correspond to the infrared active vibrational modes of the fullerene species C_{60} and little or no emission is present from Polycyclic Aromatic Hydrocarbons (PAHs). The strength of the fullerene bands in the three planetary nebulae is very similar, while the ratio of the [Ne iii]15.5μm/[Ne ii]12.8μm fine structure lines, an indicator of the strength of the radiation field, is markedly different. This raises questions about their excitation mechanism and we compare the fullerene emission to fluorescent and thermal models. In addition, the spectra show other interesting and common features, most notably in the 6–9 μm region, where a broad plateau with substructure dominates the emission. These features have previously been associated with mixtures of aromatic/aliphatic hydrocarbon solids. We hypothesize on the origin of this band, which is likely related to the fullerene formation mechanism, and compare it with modeled Hydrogenated Amorphous Carbon that present emission in this region.

Accepted for publication in ApJ Available from arXiv:1207.5794

Submillimeter line emission from LMC 30 Dor: The impact of a starburst on a low metallicity environment

Jorge L. Pineda^{1}, N. Mizuno^{2}, M. Roellig^{3}, J. Stutzki^{3}, C. Kramer^{4}, U. Klein^{5}, M. Rubio^{6}, A. Kawamura^{2}, T. Minamidani^{7}, A. Benz^{8}, M. Burton^{9}, Y. Fukui^{10}, B.-C. Koo^{11} and T. Onishi^{12}

1Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Drive, Pasadena, CA 91109-8099, USA 2ALMA-J Project Office, National Astronomical Observatory of Japan, 2-21-1 Osawa, Mitaka, Tokyo 181-8588, Japan 3KOSMA, I. Physikalisches Institut, Universität zu Köln, Zülpicher Straße 77, D-50937 Köln, Germany 4Instituto Radioastronomía Milimétrica, Av. Divina Pastora 7, Nucleo Central, 18012 Granada, Spain 5Argelander-Institut für Astronomie, Auf dem Hügel 71, D-53121 Bonn, Germany 6Departamento de Astronomía, Universidad de Chile, Casilla 36-D, Santiago, Chile 7Department of Physics, Faculty of Science, Hokkaido University, N10W8, Kita-ku, Sapporo 060-0810, Japan 8Institute of Astronomy, ETH Zürich, Switzerland 9School of Physics, UNSW, Sydney, NSW 2052, Australia 10Department of Astrophysics, Nagoya University, Chikusa-ku, Nagoya 464-8602, Japan 11Seoul National University, Seoul 151-742, Korea 12Department of Astrophysics, Osaka Prefecture University, 1-1 Gakuen-cho, Osaka 599-8531, Japan

The 30 Dor region in the Large Magellanic Cloud (LMC) is the most vigorous star-forming region in the Local Group. Star formation in this region is taking place in low-metallicity molecular gas that is exposed to an extreme farultraviolet (FUV) radiation field powered by the massive compact star cluster R 136. 30 Dor is therefore ideally suited to study the conditions in which stars formed at earlier cosmological times. Observations of (sub)mm and far-infrared (FIR) spectral lines of the main carbon-carrying species, CO, [C i] and [C ii], which originate in the surface layers of molecular clouds illuminated by the FUV radiation of young stars, can be used to constrain the physical and chemical state of the star-forming ISM. We used the NANTEN2 telescope to obtain high-angular resolution observations of the 12CO J = 4 → 3, J = 7 → 6, and ^{13}CO J = 4 → 3 rotational lines and [C i] ^{3}P_{1}–^{3}P_{0} and ^{3}P_{2}–^{3}P_{1} fine-structure submillimeter transitions in 30 Dor-10, the brightest CO and FIR-emitting cloud at the center of the 30 Dor region. We derived the physical and chemical properties of the low-metallicity molecular gas using an excitation/radiative transfer code and found a self-consistent solution of the chemistry and thermal balance of the gas in the framework of a clumpy cloud PDR model. We compared the derived properties with those in the N 159W region, which is exposed to a more moderate far-ultraviolet radiation field compared with 30 Dor-10, but has similar metallicity. We also combined our CO detections with previously observed low-J CO transitions to derive the CO spectral-line energy distribution in 30 Dor-10 and N 159W. The separate excitation analysis of the submm CO lines and the neutral carbon fine structure lines shows that the mid−J CO and [C i]-emitting gas in the 30 Dor-10 region has a temperature of about 160 K and a H_{2} density of about 10^{4} cm^{−3}. We find that the molecular gas in 30 Dor-10 is warmer and has a lower beam filling factor compared to that of N 159W, which might be a result of the effect of a strong FUV radiation field heating and disrupting the low-metallicity molecular gas. We use a clumpy PDR model (including the [C ii] line intensity reported in the literature) to constrain the FUV intensity to about χ_{0} ≈ 3100 and an average total H density of the clump ensemble of about 10^{5} cm^{−3} in 30 Dor-10.

Accepted for publication in A&A Available from arXiv:1206.4051

Emission-line stars discovered in the UKST Hα survey of the Large Magellanic Cloud; Part 1: Hot stars

Warren Reid^{1} and Quentin Parker^{1}

1Macquarie University, Australia

We present new, accurate positions, spectral classifications, radial and rotational velocities, Hα fluxes, equivalent widths and B,V,I,R magnitudes for 579 hot emission-line stars (classes B0–F9) in the Large Magellanic Cloud (LMC) which include 469 new discoveries. Candidate emission line stars were discovered using a deep, high resolution Hα map of the central 25 deg^{2} of the LMC obtained by median stacking a dozen 2 hour Hα exposures taken with the UK Schmidt Telescope (UKST). Spectroscopic follow-up observations on the Anglo–Australian Telescope (AAT), the UKST, the Very Large Telescope (VLT), the South African Astronomical Observatory (SAAO) 1.9-m and the 2.3-m telescope at Siding Spring Observatory have established the identity of these faint sources down to magnitude R ∼ 23 for Hα (4.5 × 10^{−17} ergs cm^{−2} s^{−1} Å^{−1}). Confirmed emission-line stars have been assigned an underlying spectral classification through cross-correlation against 131 absorption line template spectra covering the range O1 to F8. We confirm 111 previously identified emission line stars and 64 previously known variable stars with spectral types hotter than F8. The majority of hot stars identified (518 stars or 89%) are class B. Of all the hot emission-line stars in classes B–F, 130 or 22% are type B[e], characterised by the presence of forbidden emission lines such as [S ii], [N ii] and [O ii]. We report on the physical location of these stars with reference to possible contamination from ambient H ii emission. Only 13 of the emission-line stars require additional high resolution spectroscopic observations in order to assign a spectroscopic classification. They have nonetheless been added to the catalogue. Along with flux calibration of the Hα emission we provide the first Hα luminosity function for selected sub-samples after correction for any possible nebula or ambient contamination. We find a moderate correlation between the intensity of Hα emission and the V magnitude of the central star based on SuperCOSMOS magnitudes and the Optical Gravitational Lensing Experiment (OGLE-II) photometry where possible. Cool stars from classes G–S, with and without strong Hα emission, will be the focus of part 2 in this series.

Accepted for publication in MNRAS Available from arXiv:1207.1950

Conference Paper

New studies of a variety of Southern pulsating B stars

C. Ulusoy^{1}, C.A. Engelbrecht^{1} and T. Gülmez^{1}

1Department of Physics, University of Johannesburg, P.O. Box 524, Auckland Park 2006, Johannesburg, South Africa

We present preliminary results of multi-colour photometry of β Cephei stars observed in the LMC and in NGC 6200. Tentative identifications of pulsation modes have been made, and a number of new B pulsators have been noted. Interesting features have also been discovered in the light curves of some of these stars.

Poster contribution, published in the 20^{th} edition of the Stellar Pulsation Conference: ”Impact of new instrumentation and new insights in stellar pulsations”, Granada, 5–9 September 2011 Available from arXiv:1206.0227

Announcement

30 Doradus: The Starburst Next Door A mini-workshop at STScI

The 30 Doradus or Tarantula Nebula region of the Large Magellanic Cloud is synonymous with many superlatives in astronomy, hosting as it does the most massive young resolved cluster (R 136), the most massive stars yet discovered, the fastest rotating O-type stars, the most massive runaway star...

Its unique accessibility to detailed study covering the electromagnetic spectrum from X-ray, UV, optical, IR through radio is reflected in the many detailed surveys of this region with facilities such as HST, Chandra, Spitzer, VLT–Flames and VISTA. While it is certainly a challenge to understand this wealth of information for what is a very complex region it is one well worth addressing since 30 Doradus may have much to teach us about more distant unresolved starbursts and super star clusters.

The aim of this mini-workshop is to review recent results and outstanding theoretical issues and to examine the extent to which our knowledge of 30 Doradus can be used to improve our understanding of distant starbursts and star clusters. It will focus on five closely related themes: formation, environment, content, evolution, relevance to Super Star Clusters and Starbursts.

Invited Speakers: Daniela Calzetti, You-Hua Chu, Paul Crowther, Chris Evans, Jay Gallagher, Remy Indebetouw, Laura Lopez, Hugues Sana, Leisa Townsley, Hans Zinnecker

SOC: Selma de Mink, Linda Smith, Karl Gordon, Nolan Walborn, Danny Lennon (Chair), Brad Whitmore, Elena Sabbi, Aida Wofford, Sherita Hanna (coordinator)

Mini-workshop Dates: Sept. 16^{th}–19^{th} 2012 Registration Dead-line: August 17^{th} 2012

See also http://www.stsci.edu/institute/conference/doradus